1.50 mole each of H2 and I2 were placed in a sealed 10L container maintained at 717k .at equilibrium 1.25 mol each of H2,I2 were left behind.equilibrium constant for H2+I2=2HI

Grade:12th pass

1 Answers

Arun
5 years ago
Initial miles of H2 and I2 = 1.5 moles Volume of container=10 L At equilibrium moles of H2 and I2 left = 1.25 No.of moles H2 and I2 used = 1.5- 1.25=0.25 (each) So HI produced= 0.50 moles Kc =(0.50/10)²/(1.25/10)² Kc = 0.16